# Cosmography with next-generation gravitational wave detectors

### v1

@@ 2.4 Love sirens
The ability to generate EM counterparts and the constrained mass spectrum make neutron star mergers effective bright and spectral siren candidates, respectively. But that is not all— neutron stars that are part of compact binary coalescence (in either a BNS or an NSBH) undergo tidal deformation during the later stages of the inspiral. These deformations, at leading order, affect the phase evolution at the fifth post-Newtonian order (i.e, they are suppressed by a factor of $(v/c)^{10}$ compared to the dominant phase term, where $v$ is the orbital speed and $c$ is the speed of light) [136, 137]. The leading order tidal contribution is a linear function of the individual tidal deformability parameters of the two neutron stars, $\Lambda_{1}$ and $\Lambda_{2}$, and is called the reduced tidal deformability parameter $(\widetilde{\Lambda})$. If the neutron stars are spinning, the individual $\Lambda$s will also appear in the contributions to the phase from spin-induced quadrupolar deformations [138, 139, 140, 141]. These changes in the phase allow the measurement of $\Lambda_{1}$ and $\Lambda_{2}$ when GWs are detected. Now, given an equation of state (EOS) for neutron stars, one can relate the individual $\Lambda$s to source-frame masses of the corresponding neutron stars. These source-frame estimates, along with the detector-frame redshifted masses estimated from the GW signal, can be used to ascertain the redshift associated with the binary [142]. The post-merger signal of the BNS merger can also be used to break the redshift-mass degeneracy, by utilizing the robust spectral features that depend on the source-frame masses of the binary [143]. These approaches that use the tidal deformability measurement and spectral features of the BNS system and the knowledge of the neutron star EOS to measure the redshift and obtain bounds on cosmological parameters constitute the Love siren approach (as an ode to the neutron star tidal Love number [144]).
Note that, in practice, only $\widetilde{\Lambda}$ and not the individual $\Lambda$ are expected to be precisely measured, even with XG observatories [145]. However, there exist phenomenological relations, called quasi-universal relations [146], that are independent of the neutron star EOS and allow the estimation of individual $\Lambda$s using $\widetilde{\Lambda}$ and the mass-ratio of the binary. While this approach can indeed provide narrower bounds on the estimates of $\Lambda_{1}$ and $\Lambda_{2}$, they come at the cost of incurring biases due to uncertainties in the quasi-universal relations [146, 147].
@@
The Love siren method provides a GW-only approach to the measurement of the cosmological parameters. It relies on the knowledge of the neutron star EOS, which is expected to be well determined by the XG era [151, 152, 8]. So, either a previously inferred EOS can be used directly, or the EOS and the cosmological parameters can be jointly inferred for the GW data [153]. The forecasts that involve the post-merger spectrum require additional care, as they would strongly depend not only on the number of high-SNR BNS detections, but also on the post-merger modeling of the hypermassive neutron star remnant. Another aspect that requires caution, as noted earlier, is the use of quasi-universal relations to obtain individual $\Lambda$s from $\tilde{\Lambda}$ and the mass ratio. The fits for the universal relations themselves have uncertainties that can propagate to inference of cosmological parameters, inducing biases, especially at high precision. These uncertainties can be addressed by marginalizing over the residuals [154, 155, 156] or by specifically correcting for the incurred errors [147].
@@ 3.1 GW astronomy in 2035+
The detection of GW from compact binary mergers has opened a new window to the universe. The $\sim 100$ events detected by the end of the LVK third observing run have resulted in constraints on deviations from general relativity [157], inference of cosmological parameters [98], revelations about neutron star properties [24, 158, 159, 160, 161, 162] and measurement of the population properties of compact binary mergers [163]. However, based on the current estimates of the local merger rates and the redshift distribution of compact binary mergers, more than a million such events are expected to occur every year [8], most of which are missed by the GW networks at current sensitivities. Of the $\sim 100$ events that are detected, the majority lie close to the SNR threshold of $10$, which does not allow the precise measurement of cosmological parameters, deviations from general relativity, or neutron star properties. Thus, to realize the full potential of GW, bigger and more sensitive detector networks are needed.
@@
The first such advancement, called the A+ concept [164, 165], is planned for the second half of 2020s. With lower quantum noise and thermal coating noise, A+ is expected to be a $50\%$ improvement over advanced LIGO design sensitivity. In the most optimistic scenario, all five detectors, including the three LIGO detectors in Livingston, Hanford and the planned detector in Aundha (India), the Virgo detector in Italy, and the KAGRA detector in Japan, will have sensitivities similar to A+ by the end of 2020s. Apart from the improved sensitivity, such a 5-detector network would also excel in accurately pinpointing the location of mergers in the sky [166, 167] and play an instrumental role in dark and bright siren cosmology [64, 101, 8]. Beyond A+, one of the proposed improvements to the three LIGO detectors is the ${}^{\sharp}$ proposal [168] for the early 2030s. ${}^{\sharp}$ is expected to be about two times as sensitive as A+ across the frequency band. The improved sensitivity has a direct impact on the number of sources detected as well as the estimation of parameters, leading to more accurate distance and localization estimation, which will result in stronger constraints on cosmological parameters [8].
@@
Even after the improvement in sensitivity, the science that can be explored with the A+ and the ${}^{\sharp}$ networks is limited by the size of the detectors. This is remedied by the proposed CE and ET observatories. The CE project [169, 170, 2, 171] involves the construction of two L-shaped detectors with arms of length 40 km for one detector and 20 km for the other. The order of magnitude increase in size and noise mitigation strategies result in $\mathcal{O}(10)-\mathcal{O}(100)$ improvement in the sensitivity compared to A+. On the other hand, ET [172, 173] is the proposed underground observatory in Europe, currently planned with three detectors with arms of length 10 km placed along the vertices of an equilateral triangle in a xylophone design (a comparison of different design proposals can be found in Ref. [174], including the possibility of two underground L-shape detectors at different locations). ET is expected to have better sensitivity than CE at frequencies less than 10 Hz. Due to this, the inspiralling binaries remain in the sensitive frequency band of the observatory for much longer, which leads to an improvement in parameter estimation. Both the CE and the ET observatories are projected to begin observations in the late 2030s.
@@
Apart from the ground-based observatories discussed above, the Laser Interferometer Space Antenna (LISA) [176], a space-based GW observatory, is planned to launch in the mid-2030s. This will consist of three satellites 2.5 million kms apart, located at the vertices of an equilateral triangle, moving in a heliocentric orbit. This will allow LISA to observe GW in the mHz regime. While LISA will be able to constrain cosmological parameters to unprecedented precision by itself, the GW sources utilized for this science goal are distinct from the ones used by ground-based detectors [3]. Instead of binaries containing neutron stars, massive BBH mergers are expected to serve as bright sirens. This is because the matter accreting on the massive black holes is expected to generate EM radiation close to merger [177, 178]. Despite the low expected detection rate, the exquisite sky localization [179] and the emission of EM radiation across all wavelengths [180] render these sources as promising bright siren candidates. For dark siren measurements, stellar-mass BBHs within the redshift of $z=0.1$ can be employed as dark sirens, constraining $H_{0}$ to $\mathcal{O}(1)\%$ [181]. Extreme-mass ratio at higher redshifs $(z\sim 0.5)$ can also be used as dark sirens and can constrain $H_{0}$ to $\mathcal{O}(1)\%$ as well [182]. Moreover, as LISA can detect massive BBH sources to high redshifts $(z>1)$, it can utilize the effects of weak lensing to infer clustering parameters like $\sigma_{8}$ and $n_{S}$ [183]. These sources can also be used to test $\Lambda$CDM cosmology and probe the nature of dark energy [3]. LISA is also well positioned to test the validity of general relativity at cosmological scales [184].
@@ 3.2 EM observatories in 2035+
Corresponding EM facilities will be critical for the GW siren measurements in the XG era. The observations of kilonova AT 2019gfo enabled the association to the host galaxy of BNS GW170817, allowing for the first bright siren measurement [26, 14]. Dark siren measurements with compact binary mergers observed by LVK also relied on the availability and completeness of galaxy catalogue [115, 96, 188].
@@
Figure 4: Planned and proposed wide-field EM facilities that could contribute to the search for EM counterparts in the XG era [191, 192, 193, 194, 195, 196, 197]. The ${}^{\sharp}$  network includes the two current LIGO sites, Hanford and Livingston in the US, and the planned Aundha in India.
@@ 3.3 Cosmology landscape in 2035+
The study of the cosmos has seen a revolution in the last three decades with the discovery of the accelerated expansion of the Universe [10, 11] and the precise mapping of the anisotropies of the cosmic microwave background (CMB) [198, 199]. These new cosmological probes have established a surprising “standard” picture, called $\Lambda$CDM model, in which most of the content of the present Universe is in the form of an unknown dark energy and dark matter. Much of the efforts of the large facilities that have been built and are planned to be built is precisely the unveiling of the physical origin of the dark Universe.
@@
The main cosmological probes are the observation of the primordial Universe with the CMB, the measurement of the expansion rate with supernova type Ia (SNIa) and baryon acoustic oscillations (BAO), and the study of the growth of structure with galaxy surveys mapping the large scale structure (LSS), although there are many other emergent probes [200]. All these probes aim at stress-testing $\Lambda$CDM in order to consolidate its position or discover new physics. Several tensions have been claimed over the years, but undoubtedly the most advertised one is the “Hubble tension” [201, 202], reporting a disagreement between the measurement of the local expansion rate and the inferred value of $H_{0}$ from the CMB. It is unclear how long this tension will last, but, as we will see later, GW standard siren cosmology will provide a new way of precisely measuring the Universe’s expansion rate.
@@
There are currently multiple facilities with the potential to shed light on our understanding of dark energy and dark matter. Most notably, the Dark Energy Spectroscopic Instrument (DESI) [87] and the Euclid satellite [190] will be able to map millions of galaxies over 10 billion years of cosmic history ($z\sim 0$–2) to constrain the properties of dark energy and the growth rate of the LSS using BAO, weak lensing maps and redshift space distortions among other methods. For example, DESI is expected to constrain the expansion rate $H(z)$ to a few percent precision over this wide range of redshifts [87]. These surveys build up on the efforts of their predecessors, the Dark Energy Survey (DES) [203] and the Sloan Digital Sky Survey (SDSS) [204]. In the near future, the Vera Rubin Observatory (VRO) [205] will change the scale in which we map the Universe and observe its transient signals, playing a capital role in the multi-messenger transient endeavor. For instance, VRO is expected to detect over 10 million supernovae during a 10-year period, which can be compared with the order 1500 that DES accumulated in 5 years [206]. The Nancy Grace Roman Space Telescope [196] is expected to have a similar role from space, providing a detailed wide-field view of the sky. Roman seeks to measure both expansion history and structure growth parameters with 0.1 – 0.5% precision level, reaching 0.9% and 2.1% constraints on $H(z)$ between $z=$1-2 and $z=$2-3 respectively [196]. More into the future, the Square Kilometer Array (SKA) [207] will look further back on time to explore the Universe when only the first galaxies were starting to be formed.
@@ 4 New horizons with XG detectors
In this section, we will make projections on the measurement of cosmological parameters using GW observations. From the same, we have considered various detector networks, starting from ${}^{\sharp}$ sensitivity to a network with ET and two CE observatories. Details about the network configurations can be found in A. Specifically, we simulate populations of BBH, NSBH and BNS mergers consistent with current observations and following common astrophysical assumptions (details are described in B.1 and B.2). The golden dark siren approach has been applied to the three compact binary classes and is presented in section 4.1. In section 4.2, we generate kilonova light curves for BNS mergers and project the number of bright sirens based on the number of kilonovae that can be detected by LSST. The bounds on cosmological parameters using the spectral siren method with BBH and BNS mergers have been discussed in section 4.3. The joint forecast based on the estimates from these three sections has been discussed in section 4.4.
@@ 4.1 Golden dark sirens
Note that not all realizations of the universe will have golden dark sirens. Thus, whenever a realization has no dark siren, we simply get back the prior that was put on $H_{0}$. From figure 5, we see that the number of golden dark sirens is much less compared to the bright sirens, but the constraints on $H_{0}$ are comparable for most cases. For the ${}^{\sharp}$ network, BBH golden dark sirens can contain $H_{0}$ to $3\%$. Most of the realizations of the universe do not contain any NSBH or BNS golden dark sirens in the ${}^{\sharp}$ detections, leading to the median value coinciding with the $H_{0}$ prior. This is also the case with a network that contains just the ET observatory. However, here, even the BBH golden dark sirens are not expected to contribute significantly to $H_{0}$. This is due to the stringent sky-localization constraint on the selection of golden dark sirens, where a network of three ${}^{\sharp}$ detectors located far away from each other is seen to perform better than the three co-located detectors that make up the ET observatory. A similar observation is made for CE4020, where the two-detector network, with both detectors located in the US, is not able to measure the sky area well enough to qualify the detections as golden dark sirens. The CE40ET network does a much better job at constraining $H_{0}$, doing so to sub-percent precision for all the three classes of golden dark sirens. The constraints improve by a factor of 2-3 with the inclusion of CE20 in this network. Remarkably, for CE4020ET, all the realizations contain at least one golden dark siren, and almost all of them can constrain $H_{0}$ to better than $2\%$, resolving the $H_{0}$ tension in a year of observation.

### v2

@@ 2.4 Love sirens
The ability to generate EM counterparts and the constrained mass spectrum make neutron star mergers effective bright and spectral siren candidates. But that is not all— neutron stars that are part of compact binary coalescence (in either a BNS or an NSBH) undergo tidal deformation during the later stages of the inspiral. These deformations, at leading order, affect the phase evolution at the fifth post-Newtonian order (i.e, they are suppressed by a factor of $(v/c)^{10}$ compared to the dominant phase term, where $v$ is the orbital speed and $c$ is the speed of light) [157, 158]. The leading order tidal contribution is a linear function of the individual tidal deformability parameters of the two neutron stars, $\Lambda_{1}$ and $\Lambda_{2}$, and is called the reduced tidal deformability parameter $(\widetilde{\Lambda})$. If the neutron stars are spinning, the individual $\Lambda$s will also appear in the contributions to the phase from spin-induced quadrupolar deformations [159, 160, 161, 162]. These changes in the phase allow the measurement of $\Lambda_{1}$ and $\Lambda_{2}$ when GWs are detected. Now, given an equation of state (EOS) for neutron stars, one can relate the individual $\Lambda$s to source-frame masses of the corresponding neutron stars. These source-frame estimates, along with the detector-frame redshifted masses estimated from the GW signal, can be used to ascertain the redshift associated with the binary [163]. The post-merger signal of the BNS merger can also be used to break the redshift-mass degeneracy, by utilizing the robust spectral features that depend on the source-frame masses of the binary [164]. These approaches that use the tidal deformability measurement and spectral features of the BNS system and the knowledge of the neutron star EOS to measure the redshift and obtain bounds on cosmological parameters constitute the Love siren approach (as an ode to the neutron star tidal Love number [165]).
Note that, in practice, only $\widetilde{\Lambda}$ and not the individual $\Lambda$ are expected to be precisely measured, even with XG observatories [166]. However, there exist phenomenological relations, called quasi-universal relations [167], that are independent of the EOS governing the inspiralling neutron stars and allow the estimation of individual $\Lambda$s using $\widetilde{\Lambda}$ and the mass-ratio of the binary. While this approach can indeed provide narrower bounds on the estimates of $\Lambda_{1}$ and $\Lambda_{2}$, they come at the cost of incurring biases due to uncertainties in the quasi-universal relations [167, 168].
@@
The Love siren method provides a GW-only approach to the measurement of the cosmological parameters. It relies on the knowledge of the neutron star EOS, which is expected to be well determined by the XG era [172, 173, 7, 174]. So, either a previously inferred EOS can be used directly, or the EOS and the cosmological parameters can be jointly inferred for the GW data [175]. The forecasts that involve the post-merger spectrum require additional care, as they would strongly depend not only on the number of high-SNR BNS detections, but also on the post-merger modeling of the hypermassive neutron star remnant. Another aspect that requires caution, as noted earlier, is the use of quasi-universal relations to obtain individual $\Lambda$s from $\tilde{\Lambda}$ and the mass ratio. The fits for the universal relations themselves have uncertainties that can propagate to inference of cosmological parameters, inducing biases, especially at high precision. These uncertainties can be addressed by marginalizing over the residuals [176, 177, 178] or by specifically correcting for the incurred errors [168].
@@ 3.1 GW astronomy in 2035+
The detection of GW from compact binary mergers has opened a new window to the universe. The $\sim 100$ events detected by the end of the LVK third observing run have resulted in constraints on deviations from general relativity [179], inference of cosmological parameters [109], revelations about neutron star properties [32, 180, 181, 182, 183, 184] and measurement of the population properties of compact binary mergers [185]. However, based on the current estimates of the local merger rates and the redshift distribution of compact binary mergers, more than a million such events are expected to occur every year within 600 Gpc $(z\sim 50)$ [7], most of which are missed by the GW networks at current sensitivities. Of the $\sim 100$ events that are detected, the majority lie close to the SNR threshold of $10$, which does not allow the precise measurement of cosmological parameters, deviations from general relativity, or neutron star properties. Thus, to realize the full potential of GW, bigger and more sensitive detector networks are needed.
@@
The first such advancement, called the A+ concept [4, 5], is planned for the second half of 2020s. With lower quantum noise and thermal coating noise, A+ is expected to be a $50\%$ improvement over advanced LIGO design sensitivity. In the most optimistic scenario, all five detectors, including the three LIGO detectors in Livingston, Hanford and the planned detector in Aundha (India), the Virgo detector in Italy, and the KAGRA detector in Japan, will have sensitivities similar to A+ by the end of 2020s. Apart from the improved sensitivity, such a 5-detector network would also excel in accurately pinpointing the location of mergers in the sky [110, 112] and play an instrumental role in dark and bright siren cosmology [44, 117, 7]. Beyond A+, one of the proposed improvements to the three LIGO detectors is the A proposal [6] for the early 2030s. A is expected to be about twice as sensitive as A+ across the frequency band. The improved sensitivity has a direct impact on the number of sources detected as well as the estimation of parameters, leading to more accurate distance and localization estimation, which will result in stronger constraints on cosmological parameters [7].
@@
Even after the improvement in sensitivity, the science that can be explored with the A+ and the A networks is limited by the size of the detectors. This is remedied by the proposed CE and ET observatories. The CE project [186, 187, 8, 188] involves the construction of two L-shaped detectors with arms of length 40 km for one detector and 20 km for the other. The order of magnitude increase in size and noise mitigation strategies result in $\mathcal{O}(10)-\mathcal{O}(100)$ improvement in the sensitivity compared to A+. On the other hand, ET [189, 190] is the proposed underground observatory in Europe, currently planned with three detectors with arms of length 10 km placed along the vertices of an equilateral triangle in a xylophone design (a comparison of different design proposals can be found in Ref. [113], including the possibility of two underground L-shape detectors at different locations). ET is expected to have better sensitivity than CE at frequencies less than 10 Hz. Due to this, the inspiralling binaries remain in the sensitive frequency band of the observatory for much longer, which leads to an improvement in parameter estimation. Both the CE and the ET observatories are projected to begin observations in the late 2030s.
@@
Apart from the ground-based observatories discussed above, the Laser Interferometer Space Antenna (LISA) [192], a space-based GW observatory, is planned to launch in the mid-2030s. This will consist of three satellites 2.5 million kms apart, located at the vertices of an equilateral triangle, moving in a heliocentric orbit. This will allow LISA to observe GW in the mHz regime. While LISA will be able to constrain cosmological parameters to unprecedented precision by itself, the GW sources utilized for this science goal are distinct from the ones used by ground-based detectors [11]. Instead of binaries containing neutron stars, massive BBH mergers are expected to serve as bright sirens. This is because the matter accreting on the massive black holes is expected to generate EM radiation close to merger [193, 194]. Despite the low expected detection rate, the exquisite sky localization [195] and the emission of EM radiation across all wavelengths [196] render these sources as promising bright siren candidates. For dark siren measurements, stellar-mass BBHs within the redshift of $z=0.1$ can be employed as dark sirens, constraining $H_{0}$ to $\mathcal{O}(1)\%$ [197]. Extreme-mass ratio inspirals at higher redshifs $(z\sim 0.5)$ can also be used as dark sirens and can constrain $H_{0}$ to $\mathcal{O}(1)\%$ as well [198]. Moreover, as LISA can detect massive BBH sources to high redshifts $(z>1)$, it can utilize the effects of weak lensing to infer clustering parameters [199]. These sources can also be used to test $\Lambda$CDM cosmology and probe the nature of dark energy [11]. LISA is also well positioned to test the validity of general relativity at cosmological scales [200].
@@ 3.2 EM observatories in 2035+
Corresponding EM facilities will be critical for the GW siren measurements in the XG era. The observations of kilonova AT 2019gfo enabled the association to the host galaxy of the BNS GW170817, allowing for the first bright siren measurement [34, 20]. Dark siren measurements with compact binary mergers observed by LVK also relied on the availability and completeness of galaxy catalogue [136, 103, 106].
@@
Figure 4: Planned and proposed wide-field EM facilities that could contribute to the search for EM counterparts in the XG era [206, 207, 208, 209, 210, 211, 212, 213, 214]. The A  network includes the two current LIGO sites, Hanford and Livingston in the US, and the planned Aundha in India.
@@ 3.3 Cosmology landscape in 2035+
The study of the cosmos has seen a revolution in the last three decades with the discovery of the accelerated expansion of the Universe [16, 17] and the precise mapping of the anisotropies of the cosmic microwave background (CMB) [215, 26]. These new cosmological probes have established a surprising “standard” picture, called the $\Lambda$CDM model, in which most of the content of the present Universe is in the form of an unknown dark energy and dark matter. Much of the efforts of the large facilities that have been built and are planned to be built is precisely the unveiling of the physical origin of the dark Universe.
@@
The main cosmological probes are the observation of the primordial Universe with the CMB [215, 26], the measurement of the expansion rate with supernova type Ia (SNIa) [16, 17] and baryon acoustic oscillations (BAO) [216], and the study of the growth of structure with galaxy surveys mapping the large scale structure (LSS) [217], although there are many other emergent probes [218]. All these probes aim at stress-testing $\Lambda$CDM in order to consolidate its position or discover new physics. Several tensions have been claimed over the years, but undoubtedly the most advertised one is the “Hubble tension” [219, 220], reporting a disagreement between the measurement of the local expansion rate and the inferred value of $H_{0}$ from the CMB. It is unclear how long this tension will last, but, as we will see later, GW standard siren cosmology will provide a new way of precisely measuring the Universe’s expansion rate.
@@
There are currently multiple facilities with the potential to shed light on our understanding of dark energy and dark matter. Most notably, the Dark Energy Spectroscopic Instrument (DESI) [95] and the Euclid satellite [205] will be able to map millions of galaxies over 10 billion years of cosmic history ($z\sim 0$–2) to constrain the properties of dark energy and the growth rate of the LSS using BAO, weak lensing maps and redshift space distortions among other methods. For example, DESI is expected to constrain the expansion rate $H(z)$ to a few percent precision over this wide range of redshifts [95]. These surveys build up on the efforts of their predecessors, the Dark Energy Survey (DES) [217] and the Sloan Digital Sky Survey (SDSS) [216]. In the near future, the Vera Rubin Observatory (VRO) [221] will change the scale in which we map the Universe and observe its transient signals, playing a capital role in the multi-messenger transient endeavor. For instance, VRO is expected to detect over 10 million supernovae during a 10-year period, which can be compared with the order 1500 that DES accumulated in 5 years [222]. The Nancy Grace Roman Space Telescope [211] is expected to have a similar role from space, providing a detailed wide-field view of the sky. Roman seeks to measure both expansion history and structure growth parameters with 0.1 – 0.5% precision level, reaching 0.9% and 2.1% constraints on $H(z)$ between $z=$1-2 and $z=$2-3 respectively [211]. More into the future, the Square Kilometer Array (SKA) [223] will look further back on time to explore the Universe when only the first galaxies were starting to be formed. There are also proposals for next-generation spectroscopic surveys to follow DESI aiming at collecting more than an order of magnitude more galaxy redshifts [224].
@@ 4 New horizons with XG detectors
In this section, we will make projections on the measurement of cosmological parameters using GW observations. From the same, we have considered various detector networks, starting from A sensitivity to a network with ET and two CE observatories. Details about the network configurations can be found in A. Specifically, we simulate populations of BBH, NSBH and BNS mergers consistent with current observations and following common astrophysical assumptions (details are described in B.1 and B.2). The golden dark siren approach has been applied to the three compact binary classes and is presented in section 4.1. In section 4.2, we generate kilonova light curves for BNS mergers and project the number of bright sirens based on the number of kilonovae that can be detected by LSST. The bounds on cosmological parameters using the spectral siren method with BBH and BNS mergers have been discussed in section 4.3. The joint forecast based on the estimates from these three sections has been discussed in section 4.4.
@@ 4.1 Golden dark sirens
Note that not all realizations of the universe will have golden dark sirens. Thus, whenever a realization has no dark siren, we simply get back the prior that was put on $H_{0}$. From figure 5, we see that the number of golden dark sirens is much less compared to the bright sirens, but the constraints on $H_{0}$ are comparable for most cases. For the A network, BBH golden dark sirens can constrain $H_{0}$ to $3\%$. Most of the realizations of the universe do not contain any NSBH or BNS golden dark sirens in the A detections, leading to the median value coinciding with the $H_{0}$ prior. This is also the case with a network that contains just the ET observatory. However, here, even the BBH golden dark sirens are not expected to contribute significantly to $H_{0}$. This is due to the stringent sky-localization constraint on the selection of golden dark sirens, where a network of three A detectors located far away from each other is seen to perform better than the three co-located detectors that make up the ET observatory. A similar observation is made for CE4020, where the two-detector network, with both detectors located in the US, is not able to measure the sky area well enough to qualify the detections as golden dark sirens. The CE40ET network does a much better job at constraining $H_{0}$, doing so to sub-percent precision for all the three classes of golden dark sirens. The constraints improve by a factor of 2-3 with the inclusion of CE20 in this network. Remarkably, for CE4020ET, all the realizations contain at least one golden dark siren, and almost all of them can constrain $H_{0}$ to better than $2\%$, resolving the $H_{0}$ tension in a year of observation.